OpenStudy (anonymous):

A quiz consists of three multiple choice questions. Each question has four possible answers. Assuming equally outcomes for all answers, there are 64 possible arrangements of answers. What is the probability of getting a perfect score? The probability of all your answers being the same that is all (A's,B's, C's, D's)

OpenStudy (kropot72):

Only one of the 64 possible arrangements of answers will give a perfect score. Therefore the probability of getting a perfect score is 1/64. The probability of all A's 1/64 as is the probability of all B's and of all C's. Therefore the required probability is given by: \[\frac{1}{64} + \frac{1}{64}+\frac{1}{64}=you\ can\ calculate\]

OpenStudy (kropot72):

@onmypagrind Do you understand?

OpenStudy (anonymous):

yea i do couldnt you just do 3/64